# Long Context and Long Form Output Dataset and Benchmarking

### v2

@@ Abstract
GPT-4o-mini achieves best human scores among automatic systems in both summarization and summary expansion tasks ($\approx$ +10% and +20%, respectively). It even surpasses human output quality in the case of short summaries ($\approx$ +7%).
Overall automatic metrics achieve low correlations with human evaluation scores ($\approx$ 0.4) but moderate correlation on specific evaluation aspects such as fluency and attribution ($\approx$ 0.6).
The LCFO benchmark offers a standardized platform for evaluating summarization and summary expansion performance, as well as corresponding automatic metrics, thereby providing an important evaluation framework to advance generative AI.
@@ 1 Introduction
Robust long text generation capabilities are required to meet user demand for extensive content creation, including story writing and essay composition (Xie and Riedl, 2024), which is why recent models such as GPT-4 (OpenAI et al., 2024) are expanding the output lengths from 4k tokens in GPT-4o to 64k in the latest versions.
Figure 1: LCFO (left): annotations consist of 3 gradual summaries plus QA on each input document plus human evaluation annotations for gradual summarization and summary expansion; with these annotations, LCFO can be used for tasks such as gradual summarization, reading comprehension, summary expansion, and automatic metric evaluation. LCFO methodology (right) consists of 6 big steps.
However, evaluating the performance of Large Language Models (LLMs) on summarization and summary expansion tasks (see definitions in Section 2.1) is particularly challenging, especially when it comes to summarizing very long input documents and generating either long summaries or long summary expansions. Although there is a lot of work and interest in studying summarization evaluation (e.g. Zhang et al., 2024a), the evaluation of long text outputs is an emerging area (Que et al., 2024). Indeed, long-context input processing tasks (such as summarization or comprehension question answering applied to long documents) and long-form output production both involve high cognitive loads for humans. This may be why evaluation work in these areas may not be as mature as in others.
The dataset presented here is an exclusively human annotated benchmark and challenge dataset involving natural language understanding and generation across multiple domains in the aforementioned tasks. Our data set is carefully manually crafted with human revision from the selection of the documents to the final annotation, without relying on LLMs at any point. We provide detailed linguistic guidelines and abstractive QA. Furthermore, we provide another set of linguistic guidelines to evaluate the tasks of summarization and summary expansion.
Together, we present the LCFO benchmark. Given a source of long structured documents, we generate multiple long outputs and associated QA pairs, and we evaluate human and model outputs with human annotations. The main contributions of this work are described below (see Figure 1-left for a schematic representation of the benchmark and its tasks):
@@ 2 The LCFO Creation Framework: Principles and Methods
This section describes the detailed principles and methods that underlie LCFO and is shown in Figure 1 (right).
@@ 2.1 Definitions
We define long text as text that exceeds 5k words. For reference, the attention span for reading and taking notes has been considered 10 to 15 minutes since a 1978 seminal paper (Hartley and Davies, 1978). More recent research cited in a survey paper (Bradbury, 2016) shows that attention paid to lectures declines significantly after 20 minutes. The tasks we describe here are closer to note taking than lecture attendance; therefore, we should keep the 10–15 minute reference. If we base ourselves on the reported reading average time for first-language, secondary-education level readers (125 words per minute on average), we can consider that we will reach high cognitive load at around 1.5k words, and peak cognitive load at around 2.5k words. The quality of cognitive processing then starts to decrease significantly after 2.5 k words.
@@
The input and output documents are partitioned into sections and, if necessary, nested sub-sections. Their structure is determined by task- and domain-specific guidelines. The size of the partitions is such that it makes them shorter (see examples of such structure in the data sets details).
@@
The input is a long document (defined here as 5k words or more). The summarization task can be described as the act of generating a much shorter corresponding document that includes the essential information contained in the long document and the same logical structure linking the various pieces of essential information. The summarization task presented here consists in taking long documents as inputs and generating three corresponding summaries of length that represent 20%, 10%, or 5% of the input document.
@@
The input is a short and concise document that has similar properties to those of a summary (that is, it is mainly a standalone document that abstracts from details). The summary expansion task can be described as the act of generating a much longer document that preserves the essential elements found in the corresponding short document as well as the logical structure that connects such elements. More specifically, the task presented here consists in taking summaries as inputs and generating 3 long documents of different lengths. Each of the 3 lengths is set such that an input summary represents either 5%, 10%, or 20% of its respective expanded documents. As this is a more freely generative task, an additional requirement to be taken into consideration is that of coherence (for example, the detailed information included in one generated sentence should not contradict that included in another sentence).
@@ 2.2 Data selection and preprocessing
  • The source documents are selected to be on average 5k words / document. We prefer documents with a hierarchical structure and containing relatively few numbers, which are better suited for summarization and summary expansion tasks.
@@
  • We also filter out documents that contain toxicity using the ETOX package (Costa-jussà et al., 2023) and add manual verification to ensure the high integrity of the selected documents.
@@ 2.3 Human summary and QA-pair generation
To obtain human-written summaries of long-form texts, detailed guidelines are developed, which are shown in Appendix C. All summary writers must be native English speakers and have writing or editing experience.
@@ 2.4 Automatic output and postprocessing
We prompt the models with the human guidelines with a slight adaptation to be LLM friendly. We input the entire document without paragraph splitting. To give a fair evaluation, we prompt all LLMs in the zero-shot setting. To control the length of the LLM output, we have added additional instructions with the upper and lower bounds of the permissible words. For example, to ask the model to generate a summary of the Rlength of the source text, the prompt contains "Make sure the summary has {y} words or less.....Please write at least {x} words"""" , where x and y are determined per document with respect to the length and ratio R. In practice, we see that enforcing the length of the document right before and after the content block in the prompt gives consistent results. We give details of our summary prompts in the Appendix G.
@@
We customize prompts for each domain, plus the model-specific prompt templates. Similarly to the summarization task, the prompt contains instructions on the desired range of the generated text length. In addition, each prompt has instructions to guide the model in generating content of a certain quality (consistency, coherence, and keeping the main ideas in the summary). We prompt the model with specific formats for different domains as reported in the appendix G.
@@
We add this step because we want to perform a human evaluation of the outputs that we are creating (as detailed in the next section 2.5). The task of comparing long inputs and outputs creates a high cognitive load on human evaluators. To reduce it, we provide an approximate alignment between the input paragraphs and the segments of the output, taking advantage of the assumption that a summary usually follows the structure of the source document. First, we use dynamic programming to find a monotonic alignment path between input and output sentences that would maximize the sum of cosine similarities of the SONAR embeddings (Duquenne et al., 2023) of the two sentences. An output sentence could be aligned with multiple consecutive input sentences, potentially from different paragraphs, but we assign it to a single input paragraph with which it is aligned the most frequently. In this way, each input paragraph gets aligned to a contiguous output segment (potentially empty) in a monotonic way. This alignment helps the annotators navigate the input and output documents in a joint way.
@@ 2.5 Human Evaluation
To perform human evaluation on previously generated output, we design human evaluation guidelines inspired by previous works (Clark et al., 2023; Krishna et al., 2023; Que et al., 2024) and fully reported in Appendices E and F.
@@
The generated summaries are evaluated in two tasks. In Task 1, the annotators first read the source document and the three summaries and then rate the generated text in four aspects, including attribution, coverage of the main ideas, conciseness and readability (similar to the ’checklist’ in HelloBench, Que et al. (2024)). The annotators rate the summary on a 0-4 Likert scale and finally give an overall rating on a 0-10 Likert scale. Each summary receives its own separate set of scores.
In Task 2, the annotators validate the QA sets that were previously created by human writers. For each question in the QA sets (13–15 questions and answers), the annotators are required to determine whether the content of the summary contains enough information to answer the question (i.e., the answer is directly stated, heavily implied or logically entailed in the summary). The annotators give a YES or NO to each question and answer. For each summary, the annotators validate the whole set of QA once.
@@
Both tasks 1 and 2 involve human judgment, and to reduce the bias, 3 sets of rating from random annotators are required for each generated output. The same guidance should be used for all different domains. Detailed evaluation guidelines are included in the Appendix E.
@@
We use the same format as the previous summarization evaluation tasks and integrate some of the questions from Story Plot Generator (Zhu et al., 2023) and HelloBench (Que et al., 2024). For task 1, the annotators read the source summary and the generated long-form output, and rate the generated text on six aspects, including the coverage of main core ideas, cohesion, richness in details, creativity, non-repetitiveness, and interest, then give an overall rating at the end. In task 2 they validate the QA set with the generated long-form text. Each output is evaluated separately without reference and three sets of random annotation ratings are required. Detailed evaluation guidelines are included in the Appendix F.
@@
Summaries and summary expansions are each evaluated separately. For the evaluation of generated summaries, 252 documents from all domains are used as the source to generate the summaries (with 2 documents being excluded during the process). The summaries are generated using three different models (as reported in Section 3 and chosen to represent close and open models of different sizes): GPT-4o-mini-64k (OpenAI et al., 2024), Llama-3.1-70B, and Llama-3.1-8B (et al., 2024). This results in 756 outputs and, along with 252 sets of human-written summaries, creates a dataset of 1,000 document-summary pairs for evaluation. A vendor sources 287 annotators, who are required (1) to be native speakers of English and (2) to hold a language-related degree. These annotators are selected from a pool that is different from that of the summary writers, ensuring that they have no prior knowledge of the source documents or the written summaries. Tasks are randomly assigned to annotators until every set of generated output receives three complete annotations. A limit of 10 evaluations for each model is set per annotator to mitigate biases in the results.
For the evaluation of generated summary expansions, only a subset of data (only covering domains that do not include factual information) is selected for evaluation, including SummScreen, BookSum, SQuality and FacetSum (102 source documents in total). The expansions are generated with the same models as previously (GPT-4o-mini, Llama-3.1-70B, and Llama-3.1-8B), resulting in 306 long-form outputs. Ten experienced data analysts are selected to conduct the evaluation. Similarly to the evaluation of summaries, the tasks are assigned randomly until every long form output receives 3 complete annotations.
@@ 2.6 Automatic evaluation
The summarization outputs are typically evaluated by computing the Rouge scores (Lin, 2004) with respect to a reference. However, this approach is not sufficient for at least three reasons (Schluter, 2017): It depends too much on the reference, it offers only a comparison at the surface level, and it does not explain why a summary is good or bad. Thus, we compute several other reference-free metrics. Each targets a specific property:
@@
    Repetitiveness: how much the summary repeats the same phrases. We report the number of all the word n-grams (with n from 1 to 3) in the summary, divided by the number of such unique n-grams (REP-3) (Welleck et al., 2019).
@@
    Coherence: how similar are the sentences in the generated texts to each other. COH-2 averages similarity of the neighboring-over-one sentences (in the embedding space) Parola et al. (2023) .
@@
    Overall in order to evaluate the overall quality of the text we use the following metrics:
    • the aggregated score (AVG) from the above metrics for summarization, namely, is the average score of $-RE\text{-}3,CoLA,COH\text{-}2,SH\text{-}4,SH\text{-}5$. For summary expansion, the aggregated score (AVG) is the average score of $REP\text{-}3,CoLA,COH\text{-}2$. Note that $REP\text{-}3$ is negated to make the score monotonic. Also, for the summary expansion, $REP\text{-}3$ increases the value over the length of output, so the factor $0.2$ is empirically set to normalize the value on the 20% summary expansion task.
    • we use HelloEval (HE) score (Que et al., 2024), which is an LLM-as-judge model with different checklist trained in human evaluation.
For some model-based scores (SEAHORSE), we had to feed the whole source text to a transformer model, which was neither feasible computationally with long context inputs nor made sense given the relatively short-form training data of those models. To bypass this problem, we segment sources and summaries into aligned fragments (using a modification of the alignment algorithm in Section 2.4) with at most 50 sentences on the source side and compute model-based metrics for the segment pairs. Table 1 summarizes the list of metrics.
Task Area Metric Description Reference
Sum Attribution SH-4 Seahorse-Large-Q4 score Clark et al. (2023)
SumExp Work Count WC
Sum/SumExp Overall AVG Empirical average of metrics
Sum/SumExp Overall HE HelloEval score Que et al. (2024)
Table 1: Summary of automatic metrics used in different tasks.

### v3

@@ Abstract
GPT-4o-mini achieves best human scores among automatic systems in both summarization and summary expansion tasks ($\approx$ +10% and +20%, respectively). It even surpasses human output quality in the case of short summaries ($\approx$ +7%). Overall automatic metrics achieve low correlations with human evaluation scores ($\approx$ 0.4) but moderate correlation on specific evaluation aspects such as fluency and attribution ($\approx$ 0.6).
LCFO:
Long Context and Long Form Output Dataset and Benchmarking
Marta R. Costa-jussà, Pierre Andrews, Mariano Coria Meglioli, Joy Chen, Joe Chuang, David Dale, Christophe Ropers, Alexandre Mourachko, Eduardo Sánchez, Holger Schwenk, Tuan Tran, Arina Turkatenko, Carleigh Wood FAIR, Meta {costajussa,mortimer,mfcoria,joyqchen joe.chuang,daviddale,chrisropers,alexmourachko eduardosanchez,schwenk,tuantran,arinatur,carleighwood}@meta.com
@@ 1 Introduction
Robust long text generation capabilities are required to meet user demand for extensive content creation, including story writing and essay composition (Xie and Riedl, 2024), which is why recent models such as GPT-4 (et al., 2024b) are expanding the output lengths from 4k tokens in GPT-4o to 64k in the latest versions.
Figure 1: LCFO annotation and evaluation framework for the gradual summarization (top) and summary expansion (bottom) tasks. The resulting benchmark can also be used for other tasks such as reading comprehension and automatic evaluation metric development. In gradual summarization, long documents are split into paragraphs and annotated by humans to generate 3 level of summaries, plus a questions-answers set. Additionally, different LLMs were used to construct automated summaries at corresponding levels. In summary expansion, different human summaries are prompted to LLMs to generate the output targeting the same length as the source document. The components are described in different sections (marked in parentheses)
However, evaluating the performance of Large Language Models (LLMs) on summarization and summary expansion tasks (see definitions in Section 2.1) is particularly challenging, especially when it comes to summarizing very long input documents and generating either long summaries or long summary expansions. Although there is a lot of work and interest in studying summarization evaluation (e.g. Zhang et al., 2024a), evaluation of long text outputs is an emerging area (Que et al., 2024). Indeed, long-context input processing tasks (such as summarization or comprehension question answering applied to long documents) and long-form output production both involve high cognitive loads for humans. This may be why evaluation work in these areas is not as mature as in others.
To complement the summarization task, we include an “inverted” task of summary expansion: generating a longform text based on its shorter sketch. It requires more creativity than summarization (which can be approached extractively) but imposes more constraints than open-ended language modeling. A practical application for this task could be with a writing assistant expanding the text sketch created by the human writer into a longer output that the human subsequently revises. Another potential use is “back-translation” for summarization, with a model training to summarize the synthetic expanded documents into their original short forms.
The LCFO dataset that we are presenting in this paper is an exclusively human annotated benchmark and challenge dataset involving natural language understanding and generation across multiple domains in the aforementioned tasks. Our dataset is carefully manually crafted with human revision from the selection of the documents to the final annotation, without relying on LLMs at any point. We provide detailed linguistic guidelines and abstractive QA. Furthermore, we provide another set of linguistic guidelines to evaluate the tasks of summarization and summary expansion.
Together, we present the LCFO benchmark. Given a source of long structured documents, we generate multiple long outputs and associated QA pairs, and we evaluate human and model outputs with human annotations. The main contributions of this work are the following (see Figure 1 for a schematic representation of the dataset and tasks):
@@ 2 The LCFO Creation Framework: Principles and Methods
This section describes the detailed principles and methods that underlie LCFO (Figure 1 (right)).
@@ 2.1 Definitions
We define long text as text that exceeds 5k words. For reference, the attention span for reading and taking notes has been considered 10 to 15 minutes since a 1978 seminal paper (Hartley and Davies, 1978). More recent research cited in a survey paper (Bradbury, 2016) shows that attention paid to lectures declines significantly after 20 minutes. The tasks we describe here are closer to note taking than lecture attendance; therefore, we should keep the 10–15 minute reference. If we base ourselves on the reported reading average time for first-language, secondary-education level readers (125 words per minute on average), we can consider that we will reach high cognitive load at around 1.5k words, and peak cognitive load at around 2.5k words. The quality of cognitive processing then starts to decrease significantly after.
@@
The input and output documents are partitioned into sections and, if necessary, nested sub-sections. Their structure is determined by task- and domain-specific guidelines.
@@
The input is a long document (defined here as 5k words or more). The summarization task can be described as the act of generating a much shorter corresponding document that includes the essential information contained in the long document and the same logical structure linking the various pieces of essential information. The summarization task presented here consists in taking long documents as inputs and generating three corresponding summaries of length that represent 20%, 10%, or 5% of the input document. [^1] We did not explicitly instruct the human annotators to generate the shorter summaries hierarchically, by compressing the longer ones, but this is what they apparently did most of the time. Hence the term “gradual”: our data allows gradually interpolating between various levels of details in a text.
@@
The input is a short and concise document that has similar properties to those of a summary (that is, it is mainly a standalone document that abstracts from details). The summary expansion task can be described as the act of generating a much longer document that preserves the essential elements found in the corresponding short document as well as the logical structure that connects such elements. More specifically, the task presented here consists in taking summaries as inputs and generating 3 long documents of different lengths. Each of the 3 lengths is set such that an input summary represents either 5%, 10%, or 20% of its respective expanded documents. As this is a more freely generative task, an additional requirement to be taken into consideration is that of coherence (for example, the detailed information included in one generated sentence should not contradict that included in another sentence). We prompt the model only with one single summary and specify the length of the output we want (5x, 10x…). We do not include in the prompt longer summaries or the whole document, since they may be a possible output.
@@ 2.2 Data selection and preprocessing
  • The source documents are selected to be on average 5k words / document. We prefer documents with a hierarchical structure and containing relatively few numbers [^2] We de-prioritize documents with many numbers (such as Wikipedia pages with large tables of various statistics), because our focus is less on structured data and more on natural language, and also because comparing large sets of numbers in the source and in the summary would be a difficult task for human annotators. , which are better suited for summarization and summary expansion tasks.
@@
  • We filter out documents that contain toxicity using the ETOX package (Costa-jussà et al., 2023) and add manual verification to ensure the high integrity of the selected documents.
@@ 2.3 Human summary and QA-pair generation
To obtain human-written summaries of long-form texts, detailed guidelines are developed (Appendix E). All summary writers must be native English speakers and have writing or editing experience.
@@ 2.4 Automatic output and postprocessing
We prompt the models with the human guidelines with a slight adaptation to be LLM friendly. We input the entire document without paragraph splitting. To give a fair evaluation, we prompt all LLMs in the zero-shot setting. To control the length of the LLM output, we have added additional instructions with the upper and lower bounds of the permissible words. For example, to ask the model to generate a summary of the Rlength of the source text, the prompt contains "Make sure the summary has {y} words or less.....Please write at least {x} words"""" , where x and y are determined per document with respect to the length and ratio R. In practice, we see that enforcing the length of the document right before and after the content block in the prompt gives consistent results. We give details of our summary prompts in the Appendix I.
@@
We customize prompts for each domain, plus the model-specific prompt templates. Similarly to the summarization task, the prompt contains instructions on the desired range of the generated text length. In addition, each prompt has instructions to guide the model in generating content of a certain quality (consistency, coherence, and keeping the main ideas in the summary). We prompt the model with specific formats for different domains as reported in the appendix I.
@@
We add this step to help human annotators evaluate the outputs that we are creating (as detailed in the next section 2.5). The task of comparing long inputs and outputs creates a high cognitive load on human evaluators. To reduce it, we provide an approximate alignment between the input paragraphs and the segments of the output, taking advantage of the assumption that a summary usually follows the structure of the source document. First, we use dynamic programming to find a monotonic alignment path between input and output sentences that would maximize the sum of cosine similarities of the SONAR embeddings (Duquenne et al., 2023) of the two sentences. An output sentence could be aligned with multiple consecutive input sentences, potentially from different paragraphs, but we assign it to a single input paragraph with which it is aligned the most frequently. Thus, each input paragraph gets aligned to a contiguous output segment (potentially empty) in a monotonic way. This alignment helps the annotators navigate the input and output documents jointly.
@@ 2.5 Human Evaluation
To perform human evaluation on previously generated output, we design human evaluation guidelines inspired by previous works (Clark et al., 2023; Krishna et al., 2023; Que et al., 2024) and fully reported in Appendices G and H.
@@
The generated summaries are evaluated in two tasks. In Task 1, the annotators first read the source document and the three summaries and then rate the generated text in four aspects, including attribution, coverage of the main ideas, conciseness and readability (similar to the ’checklist’ in HelloBench by Que et al. (2024)). The annotators rate the summary on a 0-4 Likert scale and finally give an overall rating on a 0-10 Likert scale. Each summary receives its own separate set of scores.
In Task 2, the annotators validate the QA sets that were previously created by human writers. For each question in the QA sets (13–15 questions and answers), the annotators are required to determine whether the content of the summary contains enough information to answer the question (i.e., the answer is directly stated, heavily implied or logically entailed in the summary). The annotators give a YES or NO to each QA pair. For each summary, the annotators validate the whole set of QA once.
@@
Both tasks 1 and 2 involve human judgment, and to reduce the bias, 3 sets of rating from random annotators are required for each generated output. The same guidance should be used for all different domains. Detailed evaluation guidelines are included in the Appendix G.
@@
We use the same format as the previous summarization evaluation tasks and integrate some of the questions from Story Plot Generator (Zhu et al., 2023) and HelloBench (Que et al., 2024). For task 1, the annotators read the source summary and the generated long-form output, rate the output on 6 aspects, including the coverage of main core ideas, cohesion, richness in details, creativity, non-repetitiveness, and interest, and give an overall rating at the end. In task 2, they validate the QA set with the generated long-form text. Each output is evaluated separately without reference, and three sets of random annotation ratings are required. Detailed evaluation guidelines are included in the Appendix H.
@@
Summaries and summary expansions are each evaluated separately. For the evaluation of generated summaries, 252 documents from all domains are used as the source to generate the summaries (with 2 documents being excluded during the process). The summaries are generated using three different models (as reported in Section 3 and chosen to represent close and open models of different sizes): GPT-4o-mini-64k (et al., 2024b), LLaMA 3.1-70B, and LLaMA 3.1-8B (et al., 2024a). This results in 756 outputs and, along with 252 sets of human-written summaries, creates a dataset of 1,000 document-summary pairs for evaluation. A vendor sources 287 annotators, who are required (1) to be native speakers of English and (2) to hold a language-related degree. These annotators are selected from a pool that is different from that of the summary writers, ensuring that they have no prior knowledge of the source documents or the written summaries. Tasks are randomly assigned to annotators until every set of generated output receives three complete annotations. A limit of 10 evaluations for each model is set per annotator to mitigate biases in the results.
For evaluation of generated summary expansions, only a subset of data is selected, [^3] We exclude domains with high density of factual information, because we believe they are less appropriate for the summary expansion task. The problem here is potential hallucination of factual information (e.g. Berlin is the capital of France), which can be detrimental for real-life use cases but is out of scope of our evaluation that does not check for extrinsic factuality but only for faithfulness to the source text. including SummScreen, BookSum, SQuality and FacetSum (102 source documents in total). The expansions are generated with the same models as previously (GPT-4o-mini, LLaMA 3.1-70B, and LLaMA 3.1-8B), resulting in 306 long-form outputs. Ten experienced data analysts are selected to conduct the evaluation. Similarly to the evaluation of summaries, the tasks are assigned randomly until every long form output receives 3 complete annotations.
@@ 2.6 Automatic evaluation
The summarization outputs are typically evaluated by computing ROUGE scores (Lin, 2004) with respect to a reference. However, this approach is not sufficient for at least three reasons (Schluter, 2017): it depends too much on the reference, it offers only a comparison at the surface level, and it does not explain why a summary is good or bad. Thus, we compute several other reference-free metrics, each targeting a specific aspect of summarization quality (over the 6 aspects introduced in the SEAHORSE dataset by Clark et al. (2023)). For each aspect, we tried to pick a metric that is focused on this specific aspect, preferring ones that are computationally transparent and that have already been implemented elsewhere:
@@
    Repetitiveness: how much the summary repeats the same phrases. We report the count of all the word n-grams ($n\in\{1,2,3\}$) in the summary, divided by the count of such unique n-grams (REP-3) (Welleck et al., 2019).
@@
    Coherence: how similar are the sentences in the generated texts to each other. COH-2 averages similarity of the neighboring-over-one sentences in the embedding space (Parola et al., 2023).
@@
    Overall In order to evaluate the overall quality of the text, we use two metrics:
    (1) For summarization, we aggregate a score (AVG) from the above metrics, namely, averaging $-REP\text{-}3,CoLA,COH\text{-}2,SH\text{-}4,SH\text{-}5$. For summary expansion, the aggregated score (AVG) is the average score of $REP\text{-}3,CoLA,COH\text{-}2$. Note that $REP\text{-}3$ is negated to make the score monotonic. Also, for the summary expansion, $REP\text{-}3$ increases the value over the length of output, so the factor $0.2$ is empirically set to normalize the value on the summary expansion task 20%.
    (2) We use HelloEval (HE) score (Que et al., 2024): an LLM-as-judge model with various checklists trained w.r.t. human evaluation.
Selection of these metrics from a larger candidate set was partially motivated by their correlations with human annotations from Clark et al. (2023), described in Appendix D. Table 5 in Appendix C summarizes the list of metrics.
For the SEAHORSE scores, we had to feed the whole source text to a transformer model, which was neither feasible computationally with long context inputs nor made sense given the relatively short-form training data of those models. To bypass this problem, we segment sources and summaries into aligned fragments (using a modification of the alignment algorithm in Section 2.4) with at most 50 sentences on the source side and compute model-based metrics for the fragment pairs.